OpenStudy (anonymous):

Juice solution A contains 30% juice. Juice solution B contains 50% juice. The two juice solutions are mixed together resulting in a 44% juice solution mixture. For any volume of juice solution mixture, what is the ratio of the part of the mixture that is juice A to the total amount of mixture? a. 3:10 b. 3:8 c. 3:7 d. 3:5

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\text{this is an easy example of alligation and mixture }\\}\end{align}\)

OpenStudy (mathmath333):



OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\text{u have to find n1/n2 here}\\}\end{align}\)

OpenStudy (anonymous):

I don't understand

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{n_1}{n_2}=\dfrac{A_2-A_w}{A_w-A_1}\\~\\ =\dfrac{50-44}{44-30}}\end{align}\)

OpenStudy (mathmath333):



OpenStudy (anonymous):

6/14

OpenStudy (anonymous):

3/7 simplified ?

OpenStudy (mathmath333):

yes i think so